OpenStudy (anonymous):

How does one solve the 2nd-order linear homogenous ODE? Thanks for any help.

OpenStudy (anonymous):

\[ \nabla ^2f(r)=0 \]Where: \[ r=|\vec r| \]

OpenStudy (anonymous):

This can be re-written as \[ \frac{d^2f}{dr^2}+\frac{2}{r}\frac{df}{dr}=0 \]

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

For reference the general answer is: \[ f(r)=\frac{c_1}{r}+c_2 \]

OpenStudy (anonymous):

Oh, never mind, I'm just absolutely not right today. Simple: distribute the \(r\), integrate by parts on the first term and you've got yourself a first-order linear ODE.